Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 268     Entry time: Tue Jul 23 13:50:43 2019
Author: Yusuke, Patricia 
Type: Problem Fixed 
Category:  
Subject: Reboot Flywheel PLC 
Status:  

Regarding the problem on the Flywheel PLC (see here), we entered UC1 under the supervision of Daniel M. and Koji (11:16 local time).

First we checked the LEDs of the black Ethernet cable. LINK in green was constantly ON and ACT in orange was blinking.

Then, we disconnected the black cable and reconnected again after 20 sec (11:30). But, the LEDs of the PLC in the right side were OFF.

Finally, we removed the two red power lines connected to the PLC with a screw driver and reconnected again (11:52). The LEDs of the PLC turned ON. The UC1 web page was back (12:08).

ELOG V3.1.5-fc6679b